How do I write an equation in standard form using only integers for the line through (1,4), parallel to y=-3x?
Found 2 solutions by josgarithmetic, TimothyLamb:
Answer by josgarithmetic(39617)

Slopes are the same if parallel.
y=-3x+b and you have the point which this line must contain. Plug in the values after solving for b.

b=y-(-3x)
b=y+3x
b=4+3*1
b=7
-
Your target equation is y=-3x+7.

Now adjust to form, as Ax+By=C.

Answer by TimothyLamb(4379)

---
given:
(1,4)
parallel to:
y = -3x
---
parallel slope = -3
---
to solve the linear equation, copy and paste this (the given point and parallel slope):
1,4,-3
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (2.33333333, 0)
y-intercept = (0, 7)
slope = -3
---
answer:
the slope-intercept form of our new perpendicular line:
y = -3x + 7
in standard form:
3x + y = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php